# GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and A Comprehensive Multimodal Dataset Towards General Medical AI

### v1

@@ Abstract
Despite significant advancements in general artificial intelligence, such as GPT-4, their effectiveness in the medical domain (general medical AI, GMAI) remains constrained due to the absence of specialized medical knowledge. To address this challenge, we present GMAI-VL-5.5M, a comprehensive multimodal medical dataset created by converting hundreds of specialized medical datasets into meticulously constructed image-text pairs. This dataset features comprehensive task coverage, diverse modalities, and high-quality image-text data. Building upon this multimodal dataset, we propose GMAI-VL, a general medical vision-language model with a progressively three-stage training strategy. This approach significantly enhances the model’s ability by integrating visual and textual information, thereby improving its ability to process multimodal data and support accurate diagnosis and clinical decision-making. Experimental evaluations demonstrate that GMAI-VL achieves state-of-the-art results across a wide range of multimodal medical tasks, such as visual question answering and medical image diagnosis. Our contributions include the development of the GMAI-VL-5.5M dataset, the introduction of the GMAI-VL model, and the establishment of new benchmarks in multiple medical domains. Code and dataset will be released at https://github.com/uni-medical/GMAI-VL.
@@ 1 Introduction
Large-scale Vision-Language Models (LVLMs) have rapidly evolved in recent years, effectively integrating visual perception with language understanding by leveraging large-scale multimodal data, which enables them to capture complex visual and textual patterns and drive significant advancements in image recognition, natural language processing, and multimodal tasks. With the advancement of multimodal integration technologies, the demand for high-precision processing of diverse data types in the medical field has become increasingly critical. The ability to effectively integrate and analyze various data modalities, such as medical images, clinical text, and structured clinical records, is pivotal for achieving accurate and comprehensive diagnostic and treatment decisions.
@@
However, existing LVLMs, such as GPT-4 [1], are limited in medical applications due to their lack of domain-specific knowledge, highlighting the need for specialized solutions that effectively integrate medical expertise. Addressing this challenge requires constructing a comprehensive medical vision-language dataset and developing domain-specific models. For the medical dataset, it should provide high-quality medical knowledge, including the following three aspects:
Comprehensive medical tasks. To enhance the model’s applicability across various medical scenarios, the dataset should cover a wide range of medical contexts, such as disease types, symptoms, and treatments. Comprehensive task coverage can improve the model’s generalization ability and increase its reliability in real-world applications. However, existing models often focus on specific domains [28, 41, 72, 65, 35], limiting their broader applicability. Expanding the dataset’s scope will further enhance the model’s utility in clinical practice.
Rich multimodal representation. A well-rounded medical multimodal dataset should encompass various modalities, including different medical imaging types (such as CT, MRI, and X-rays) and diverse forms of textual data (such as medical records and imaging reports). This would allow models to better integrate multi-source information and improve their analytical capabilities. However, existing methods tend to focus on a single type of medical imagery [33, 70, 50], limiting the model’s adaptability to diverse clinical scenarios. A more diverse multimodal dataset would provide a foundation for developing more comprehensive models, better suited to the complexity of real-world medical environments.
@@
Figure 1: Overview of GMAI-VL and GMAI-VL-5.5M. (a) illustrates the sources, departments, modalities, task types, and instruction formats of the GMAI-VL-5.5M dataset. (b) Architecture of GMAI-VL, integrating a Vision Encoder, Projector, and Large Language Model. (c) Three-stage training process of GMAI-VL, including shallow alignment, deep alignment, and instruction tuning with corresponding data sizes and training components. The flame symbol Refer to caption denotes the training part, while the snowflake symbol Refer to caption indicates frozen part.
High-quality image-text data. High-quality training data is essential for optimizing model performance, especially in medical applications. Ideally, image-text data for medical contexts should consist of a large, well-curated collection of medical images paired with precise textual descriptions. This data would enhance the model’s understanding of key medical concepts, such as diagnosis, treatment, and clinical workflows, ultimately contributing to improved clinical outcomes. While progress has been made with datasets sourced from platforms like PubMed [39, 54, 70], these datasets suffer from inconsistent data quality, imprecise alignment between images and text, and a lack of standardization, which limits their effectiveness.
Based on above observations, we propose an annotation-guided data generation methodology for developing a comprehensive multimodal medical datasets. The methodology begins by collecting large-scale, open-source medical imaging datasets and extracting key annotations, such as modality, task type, labels, and bounding boxes. A vision-language model (e.g., GPT-4o) is then used to transform these datasets, covering tasks like lesion detection, segmentation, and disease diagnosis, into high-quality image-text pairs for training LVLMs. To ensure data quality, extracted image information is incorporated into the prompt design, improving model performance across various clinical tasks. This results in a comprehensive multimodal dataset with 5.5M samples, named GMAI-VL-5.5M, which supports the development of general medical LVLMs. Fig. 1(a) illustrates the sources, departments, modalities, task types, and instruction formats of the constructed dataset.
@@
With the constructed GMAI-VL-5.5M dataset, we develop a general medical vision-language model, GMAI-VL. To enhance its integration of visual and linguistic features and its instruction-following abilities, a three-stage training strategy is proposed in this paper. Specifically, we sequentially implement shallow and deep alignments in the first two stages, gradually building associations between visual (medical images) and language (medical texts) elements from basic features to high-level semantics. Next, we fine-tune the model with cross-modal instructions, improving its understanding of visual-language interactions and instruction-following in complex tasks. With this strategy, GMAI-VL shows strong performance in medical tasks like visual question answering and medical image diagnosis, providing a solid foundation for advancing multimodal models in the medical field.
@@
  • We build an annotation-guided data generation methodology to construct GMAI-VL-5.5M, a comprehensive vision-language dataset. This dataset offers extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, providing a robust foundation for model training.
  • Leveraging GMAI-VL-5.5M, we design GMAI-VL, a versatile medical vision-language model. Our proposed three-stage training strategy enhances its ability to integrate visual and linguistic features, significantly improving its performance in instruction-following and generalization across a wide range of medical tasks.
  • GMAI-VL outperforms existing models in multimodal question-answering tasks, including PMC-VQA and VQA-RAD. It sets new benchmarks on OmniMedVQA, GMAI-MMBench, and Health & Medicine track of MMMU. Specifically, GMAI-VL achieves an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
@@ 2 Related Work
Large-scale medical vision-language datasets of high quality and multiple modalities are the basis of Large Vision-Language Models (LVLMs) in the medical domain. While natural language and vision datasets are easily accessible online, biomedical datasets often focus on text or images only and many of them are limited to specific tasks or modalities, thus with unsatisfactory generalization ability. Notable datasets like MIMIC-CXR [33] and CheXpert [12] have advanced radiology models but are restricted to single image modality (X-ray), which hinders their use as general-purpose medical LVLMs. To address this, researchers have begun scraping public sources like PubMed and textbooks to construct large-scale vision-language datasets. Examples include datasets proposed in LLaVA-Med [39], Med-Flamingo [54], and PubMedVision [14], with PubMedVision optimizing LLaVA-Med dataset for higher-quality medical data. In addition to scraping, open-source image datasets with annotations can also be converted into image-text pairs for model training. Specifically, image information like modalities and annotations are input into large language models, e.g., GPT series, to generate text paired with the corresponding image. Some popular examples include the datasets constructed in RadFM [70], MedDr [28], MedTrinity-25M [71], ChiMed-VL [48], BiomedGPT [79], Med-Gemini [58], and Med-PaLM [59].
@@
These efforts usually suffer from either limited modalities, data sources, or task coverage. Thus, their dataset quality needs further improvement. To this end, we propose to construct a comprehensive medical vision-language dataset with extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, forming a robust foundation for model training.
Medical vision-language models
are usually based on general-purpose Large Vision-Language Models (LVLMs). Most of them adapt LVLMs to specific medical applications using specialized medical datasets. For instance, Med-Flamingo [54] enhances OpenFlamingo-9B using 0.8 million interleaved and 1.6 million paired medical image-text data, highlighting the critical need for multimodal data in medical image analysis and automated report generation tasks. RadFM [70] improves PMC-LLaMA [69] by leveraging 16 million radiology images with text descriptions from diverse sources. Similarly, Med-PaLM [68] adapts PaLM-E [25] to the medical domain with approximately one million medical data samples, achieving state-of-the-art performance in diagnostic support and medical knowledge Q&A. LLaVA-Med [39] utilizes a large-scale biomedical figure-caption dataset extracted from PubMed Central to enhance LLaVA [66, 67] to better understand biomedical images and facilitate open-ended conversational interactions. Med-Gemini [58] leverages long-format question-answering datasets to improve the multimodal and long-contextual capabilities of the baseline Gemini model, enabling superior performance in complex medical Q&A and multimodal reasoning tasks. Additionally, HuatuoGPT-Vision [14] and MedDr [28] build medical multimodal datasets to adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
Previous studies usually focus on constructing medical datasets to adapt general-purpose LVLMs but pay less attention to the adaptation strategies. However, naive training/adaptation strategies may not successfully adapt general-purpose LVLMs to the medical data, due to the large gap between the natural image-text pairs and the medical ones. Moreover, these strategies can hardly align the broad imaging modalities and various types of medical text (e.g., prescriptions, radiology reports, and electronic health records) to obtain generalizable features, thus limiting the models’ performance. Our work thus proposes a novel three-stage training strategy to better integrate the visual and language features to enhance generalization ability.
3 GMAI-VL-5.5M: A Comprehensive Multimodal Dataset
With rapid advancements in medical vision-language models (VLMs), constructing high-quality datasets is crucial for developing general-purpose models. Unlike previous approaches that primarily rely on published literature, our method utilizes specialized medical datasets to create a more robust, high-quality resource. We build GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources. Covering 13 medical imaging modalities and 18 specialties, it addresses a wide range of medical imaging tasks. This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
@@
Table 1: Comparison of various medical multimodal datasets, including details on the dataset size, modality type, language, data traceability, and sources of information.
@@ 3.1 Data Curation
Data preprocess.
After data collection, we apply a preprocessing workflow to extract 2D medical images from the videos and 3D medical volumes. For data cleaning, we first extract key annotation information from classification data, including modality, department, and labels for each image, discarding instances with missing or unclear annotations. For segmentation data, we follow the SA-Med2D-20M [75] preprocessing method, filtering out low-quality images and labels, and converting them into detection datasets. The converted segmentation data is then processed together with the original detection dataset, with key annotation information extracted and linked to each image. Finally, the preprocessed data are then standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>.
Data generation.
Subsequently, the data are categorized into two primary types: classification datasets and detection/segmentation datasets. Each category is further refined using specific prompts tailored for large model training. For data generation, large vision-language models (GPT-4o) are employed to produce detailed image descriptions and corresponding instruction-following data through annotation-guided data generation methodology. For classification datasets, detailed descriptions of the entire image are generated, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, providing comprehensive functional analyses of these areas. Notably, the segmentation dataset was transformed into a detection dataset using external bounding boxes, and data generation followed detection dataset protocols. Furthermore, to improve the model’s multilingual capability, we translated a portion of English image-text data into Chinese. Incorporating multilingual data helps to enhance the generalization capabilities of domain-specific multimodal models. The resulted data is utilized for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, named GMAI-VL-5.5M. The detailed pipeline for generating prompt-driven data is illustrated in Fig. 2.
Figure 2: The prompt-driven data generation pipeline comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs. Figure with complete prompt and response is provided in Supp. Mat..
As depicted in Fig. 1(a), it contains six kinds of instrcuction-following formats, including image-level captions, region-level captions, free instructions, dialogue, visual perception and text-only tasks. The specific composition of GMAI-VL-5.5M can be found in Supp. Mat. (Table 6). These formats enable VLMs to better understand and process complex visual and textual information in medical contexts. The GMAI-VL-5.5M dataset significantly enhances the model’s cross-modal reasoning ability, enabling it to handle complex multimodal inputs in real clinical scenarios. The richness of the instruction formats allow the model to progress from basic question answering to advanced medical image analysis, ultimately providing strong support for clinical diagnosis and decision-making.
@@ 3.2 Data Property
These datasets encompass diverse medical imaging tasks and modalities, forming a solid foundation for developing and evaluating medical LVLMs. Fig. 4 (in Supp. Mat.) illustrates the distribution of modalities, tasks, clinical departments, and specific medical challenges represented within the collected datasets. This visualization highlights the extensive diversity and coverage of our data collection efforts. After careful standardization and integration, these datasets form the core of our comprehensive medical image-text dataset, GMAI-VL-5.5M, which serves as a crucial resource for advancing precision medicine and intelligent diagnostic systems.
Accuracy and reliability of generated data.
The accuracy and reliability of the generated data are primarily addressed from two key aspects. First, the reliability of the data source is ensured by utilizing high-quality datasets from reputable sources. The data used for generation primarily comes from professional medical challenges (such as Kaggle and GrandChallenge) and publicly available, peer-reviewed datasets. This guarantees the integrity and accuracy of the data at its source. Second, the data generation process is rigorously controlled to ensure accuracy and reliability. Although GPT was used for data generation, prompts were meticulously crafted to include essential annotations, minimizing errors and hallucinations. As illustrated in Fig. 2, the proposed annotation-guided approach incorporates specific annotations (e.g., <image, modality, label, department, bbox [optional]>) to produce precise, high-quality descriptions. Compared to the without-annotation-guided approach, which often results in less detailed outputs, the annotation-guided method delivers more professional, pathology-specific descriptions, significantly enhancing the reliability of the generated data.
@@
The GMAI-VL-5.5M dataset, as highlighted in Table 1, stands out due to its unmatched scale, encompassing over 5.5 million samples from more than 219 specialized medical imaging datasets. Unlike other datasets listed, GMAI-VL-5.5M supports a wider variety of modalities and languages, making it a truly global resource that caters to diverse clinical needs. Additionally, GMAI-VL-5.5M emphasizes traceability of its data, ensuring a high standard of clinical relevance and reliability. This comprehensive and diverse dataset is critical for pushing the boundaries of medical multimodal research, enabling more effective training of LVLMs that can generalize across multiple medical tasks and scenarios, thereby driving innovations in precision medicine and intelligent diagnostics.

### v2

@@ Abstract
Despite significant advancements in general AI, its effectiveness in the medical domain is limited by the lack of specialized medical knowledge. To address this, we formulate GMAI-VL-5.5M, a multimodal medical dataset created by converting hundreds of specialized medical datasets with various annotations into high-quality image-text pairs. This dataset offers comprehensive task coverage, diverse modalities, and rich image-text data. Building upon this dataset, we develop GMAI-VL, a general medical vision-language model, with a three-stage training strategy that enhances the integration of visual and textual information. This approach significantly improves the model’s ability to process multimodal data, supporting accurate diagnoses and clinical decision-making. Experiments show that GMAI-VL achieves state-of-the-art performance across various multimodal medical tasks, including visual question answering and medical image diagnosis.
@@ 1 Introduction
Recent advancements in Large-scale Vision-Language Models (LVLMs) have driven progress in image recognition, natural language processing, and multimodal tasks, leveraging the power of multimodal datasets. In the medical field (general medical AI, GMAI), as these technologies mature, the need for accurate processing of diverse data—such as medical images, clinical text, and structured records—has become critical for reliable diagnostic and treatment decisions.
@@
However, existing LVLMs, like GPT-4o [1], face limitations in medical applications due to their lack of domain-specific knowledge. This highlights the need for specialized solutions that incorporate medical expertise. To address this, we develop a comprehensive medical vision-language dataset and corresponding domain-specific models.
As shown in Fig. 1 (a), Our dataset provides high-quality medical knowledge across three aspects: (i) Comprehensive medical tasks: To improve the model’s applicability in various medical scenarios, our dataset covers a wide range of disease types, symptoms, treatments, and medical workflows. (ii) Rich multimodal representation: Our dataset includes a variety of modalities, such as different types of medical images (e.g., CT, MRI, X-rays) and textual data (e.g., medical records, imaging reports). (iii) High-quality image-text data: The performance of medical LVLMs heavily relies on high-quality image-text pairs. We use the well-curated collection of medical images paired with precise textual descriptions to build our dataset.
@@
Figure 1: Overview of GMAI-VL and GMAI-VL-5.5M. (a) Sources, departments, modalities, task types, and instruction formats of GMAI-VL-5.5M. (b) Architecture of GMAI-VL, with a Vision Encoder, Projector, and Large Language Model. (c) Three-stage training process, including shallow alignment, deep alignment, and instruction tuning, with corresponding data sizes and components. Refer to caption indicates the training part while Refer to caption indicates the frozen part.
To achieve this, we develop an annotation-guided data generation process to create this large-scale multimodal medical dataset. First, we collect multiple open-source medical imaging datasets and extract their key annotations (e.g., modality, task type, labels, bounding boxes). Next, we convert these well-labeled annotations into text descriptions suitable for training vision-language models. This ensures that the medical image-text pairs are accurately constructed, with annotations verified by medical experts in the collected medical imaging datasets. The process results in the GMAI-VL-5.5M dataset, consisting of 5.5 million samples, which supports the development of general medical LVLMs.
@@
Building upon this dataset, we develop a general medical vision-language model, GMAI-VL. To optimize its ability to integrate visual and linguistic features and follow complex instructions, we present a three-stage training strategy. The first two stages involve shallow and deep alignments, gradually establishing connections between medical images and texts, from basic features to high-level semantics. In the final stage, the model is fine-tuned with cross-modal instructions, improving its understanding of visual-language interactions and its ability to follow intricate tasks, as shown in Fig. 1 (b)&(c).
We conduct experiments on various multimodal medical benchmarks [39, 30, 44, 77, 16, 32, 75], and the results demonstrate that GMAI-VL significantly outperforms existing models. Specifically, GMAI-VL sets new benchmarks with an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
@@
  • We develop an annotation-guided data generation methodology to create GMAI-VL-5.5M, a comprehensive medical vision-language dataset with diverse medical tasks, rich multimodal representations, and high-quality image-text pairs.
  • Leveraging this dataset, we design GMAI-VL, a general medical vision-language model, and propose a three-stage training strategy to enhance its integration of visual and linguistic features, improving its performance in various medical tasks.
  • We demonstrate that GMAI-VL outperforms existing models in multimodal question-answering tasks, setting new benchmarks with high performance on OmniMedVQA, GMAI-MMBench, and the Health & Medicine track of MMMU.
@@ 2 Related Work
Large-scale medical vision-language datasets. These datasets are crucial for training Large Vision-Language Models (LVLMs) in the medical domain. While general datasets are readily available, biomedical datasets often focus on text or images separately, limiting their generalization. Datasets like MIMIC-CXR [35] and CheXpert [12] advance radiology models but are restricted to a single image modality (X-ray), limiting their use as general-purpose medical LVLMs.
To address this, researchers have compiled large-scale vision-language datasets by scraping public resources like PubMed and medical textbooks. Datasets like LLaVA-Med [40], Med-Flamingo [53], and PubMedVision [14] improve upon LLaVA-Med by enhancing the quality of medical data. Additionally, open-source image datasets with annotations have been converted into image-text pairs for training. Notable examples include RadFM [68], MedDr [29], BiomedGPT [76], Med-Gemini [57], and Med-PaLM [58]. MedTrinity-25M [69] generates image-text pairs for supervised fine-tuning.
@@
However, these datasets often face limitations in modalities, data sources, or task coverage, necessitating improvements. We construct a comprehensive medical vision-language dataset with broad task coverage, diverse modalities, and high-quality image-text pairs to provide a solid foundation for model training.
@@
Table 1: Comparison of multimodal medical datasets, including size, modality, language, traceability, and sources.
@@
Medical vision-language models.
Medical vision-language models often adapt general-purpose Large Vision-Language Models (LVLMs) for specific medical tasks using specialized datasets. For example, Med-Flamingo [53] improves OpenFlamingo-9B with 0.8 million interleaved and 1.6 million paired medical image-text data for medical image analysis and report generation. RadFM [68] enhances PMC-LLaMA [67] with 16 million radiology images and text descriptions. Med-PaLM [66] adapts PaLM-E [25] to medical data, achieving state-of-the-art results in diagnostic support and Q&A. LLaVA-Med [40] uses a biomedical figure-caption dataset from PubMed Central to improve LLaVA [64, 65] for biomedical image understanding and open-ended conversations. Med-Gemini [57] leverages long-format question-answering datasets for better multimodal and contextual capabilities, enhancing complex medical Q&A and reasoning tasks. Additionally, HuatuoGPT-Vision [14] and MedDr [29] adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
While these studies focus on constructing medical datasets, they often overlook adaptation strategies. Naive training methods may fail to bridge the gap between natural and medical image-text pairs, or align diverse medical modalities and texts (e.g., prescriptions, radiology reports, EHRs), limiting generalizability. Our work introduces a novel three-stage training strategy to better integrate visual and language features, improving generalization.
Table 2: The examples of annotation-guided prompts for paired data format generation.
Instruction-following data prompt
[Uncaptioned image] As a medical expert, you will receive a <fundus>image belonging to <Ophthalmology>, which is labeled as <moderate nonproliferative diabetic retinopathy>. Please generate question-answer pairs based on the following requirements: 1. The question-answer pairs must be strictly based on the content of the image and directly related to the label information, should not state known medical facts or general definitions directly. 2. Focus on reasoning around the label and guide observation of the visible features of the disease, such as affected areas, symptoms, and changes in the retina. 3. Each question and answer must be precise, clear, and derived from observable characteristics in the image, with no vague or uncertain information. 4. The output format must follow the standard JSON format with the structure as follows: { ’question’: <question content >, ’answer’: <answer content >}
Question Can we see any hard exudates or fluid leakage in the image?
Answer The image does show some bright, yellowish deposits, which are consistent with hard exudates, often seen in moderate nonproliferative diabetic retinopathy due to fluid leakage.
Question What is the appearance of the optic disc in this retinal image?
Answer The optic disc appears normal without any visible signs of swelling or abnormal changes, suggesting that the optic nerve is not yet affected by the disease.
Caption data prompt
[Uncaptioned image] As a medical expert, you are given a <X-ray> image belonging to <Pulmonary Medicine>, where the <bbox [179,164,500,434]> is labeled as <Pneumothorax>. Please generate a very detailed description related to bbox in the given image. When generating the description, you must strictly adhere to the following guidelines: 1. The generated description should be inferable from the image; 2. The generated description should include image features related to the given label; 3. The generated description should start with ’This region’; 4. The description may imply known information but should not directly state known information; 5. The generated description should include major entities within and around the region, background elements, texture features, and any other notable features.
Output: This region shows … the right hemithorax on a frontal chest X-ray, … an area suggestive of abnormal air accumulation. The bounded area appears less dense compared to surrounding lung tissue, … The edges of the lung are visible, separated from the chest wall by a faint radiolucent (darker) space, suggestive of pneumothorax. The lung markings are notably absent within this area, … The texture within the defined area has a homogeneous, darkened appearance consistent with air rather than tissue density, contrasting with the granular pattern in the aerated lung parenchyma outside this region.
3 GMAI-VL-5.5M: A Comprehensive Multimodal Medical Dataset
With rapid advancements in medical vision-language models, constructing high-quality datasets is essential for developing general-purpose models. Unlike previous approaches that mainly rely on published literature, our method leverages specialized medical datasets with various annotations to create a more robust, high-quality resource.
We present GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources. Covering 13 medical imaging modalities and 18 specialties, it supports a wide range of medical imaging tasks. This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
@@ 3.1 Data Curation
Data processing.
After data collection, we follow a workflow to pair medical data (including both 2D and 3D data) with corresponding text annotations. To ensure high-quality annotations, we first extract key information from the annotations provided by medical experts. For classification data, we extract the modality, department, and labels for each image, discarding instances with missing or unclear annotations. For segmentation data, we follow the SA-Med2D-20M [72] approach, filtering out low-quality images and labels, and converting them into detection annotation format. Finally, the preprocessed data is standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>, where “bbox” refers to the bounding box locations for detection annotations.
Paired data format generation.
To generate paired visual medical data with text descriptions, we use large vision-language models (i.e., GPT-4o) to produce detailed descriptions with instructions via an annotation-guided methodology. In details, for classification datasets, comprehensive descriptions of the entire image are created, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, with detailed functional analyses of these areas. Furthermore, based on the given information, instruction-following question-answer pairs related to the medical images are generated. These pairs involve specific instructions tailored to the medical context, such as identifying critical anatomical or pathological features within the medical images (e.g., tumors, lesions, or organs) or providing in-depth interpretations of regions of interest. These instructions guide the model to produce relevant, precise responses, thereby enhancing its applicability to specialized medical imaging tasks. As mentioned in the previous subsection, the segmentation dataset is transformed into a detection format using external bounding boxes, and data generation follows the detection dataset protocols. The detailed example prompts are shown in Table 2. The generated data is then used for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, GMAI-VL-5.5M. To enhance the model’s multilingual capability, we translate approximately 30% of the English image-text data into Chinese. This multilingual data helps further improve the generalization ability of domain-specific multimodal models.
@@ 3.2 Data Property
Our dataset encompasses a broad spectrum of medical imaging tasks and modalities, forming a robust foundation for the development and evaluation of medical LVLMs. Table 3 summarizes the distribution of key modalities, tasks, clinical departments, and clinical tasks within GMAI-VL-5.5M, illustrating its diversity and extensive coverage. For a more detailed analysis, additional visual insights into the dataset composition can be found in the supplementary material.
Data quality.
The quality of the generated paired data and annotations is ensured through two main approaches. First, we use high-quality datasets from trusted sources, including professional challenges like Kaggle and GrandChallenge, as well as peer-reviewed datasets, ensuring data accuracy and reliability. Second, we carefully control the data generation process. While GPT is used for generation, the prompts are designed with essential annotations (e.g., <image, modality, label, department, bbox [optional]>) to minimize errors. This annotation-guided approach produces more detailed and professional descriptions. The data quality is detailed in the supplementary material.
Table 3: Distribution of GMAI-VL-5.5M across key dimensions.
Dimension Value Percentage
Task Type 2D Classification 50.4%
3D Segmentation 30.3%
2D Segmentation 12.7%
2D Detection 6.6%
Modality CT 26.8%
MR 24.7%
Endoscopy 12.6%
Pathology 11.2%
X-Ray 6.7%
Fundus 5.3%
Ultrasound 3.0%
Department Orthopedic Surgery 12.9%
General Surgery 10.3%
Gastroenterology 9.7%
Hematology 9.2%
Pulmonary Medicine 9.0%
Sports Medicine 8.2%
Clinical Task Disease Diagnosis 40.4%
Organ Recognition 16.0%
Bone Recognition 8.5%
Severity Grading 6.1%
Surgeon Action Recognition 6.0%
@@
The GMAI-VL-5.5M dataset, as shown in Table 1, distinguishes itself with its unmatched scale, featuring over 5.5 million samples from more than 219 specialized medical imaging datasets. Unlike other datasets, GMAI-VL-5.5M supports a broader range of modalities and languages, making it a truly global resource that addresses diverse clinical needs. Furthermore, GMAI-VL-5.5M ensures data traceability, maintaining high clinical relevance and reliability. This extensive and varied dataset is crucial for advancing medical multimodal research, enabling more effective training of LVLMs that can generalize across numerous medical tasks and scenarios, ultimately driving innovations in precision medicine and intelligent diagnostics.